# Distributed computing quantum unitary evolution

### 11 Mar 2024 v1

@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology. When simulating chemical and biological reactions, a large number of particles are often involved, and the dimension of the quantum system composed of these particles increase exponentially as the number of particles increases, thus causing a problem called the curse of dimensionality — raised by R.E. Bellman [1, 2] to describe a number of events that occur while organizing and evaluating data in high-dimensional areas. The curse of dimensionality has always been a major obstacle to the study of high-dimensional quantum systems. However, the study of the structure of biological macromolecules is one of the most important frontier studies of quantum computing, thus the curse of dimensionality is an urgent issue to be solved. With the development of supercomputers in recent decades, the use of distributed computing can solve a series of memory and efficiency problems caused by the curse of dimensionality to a certain extent. Moreover, some distributed computing algorithms can be applied to quantum unitary evolution. A key contribution of this paper is the cavity quantum electrodynamics (QED) model, which is easy to implement in the laboratory and offers a unique scientific paradigm for studying light-matter interaction. According to this paradigm, impurity two-level systems, also known as atoms, are connected to fields of cavities. The cavity QED model includes the Jaynes–Cummings model (JCM) [3] and the Tavis-Cummings model (TCM) [4] as well as their generalizations. Many studies have been conducted recently in the field of these models [5, 6, 7, 8, 9]. In this paper, we try to simulate the cavity QED model with a large number of two-level artificial atoms on a supercomputing platform to reveal the rules of unitary evolution of a complex quantum system, and study the time cost and speedup under different distributed computing strategies comparing with that under the situation without distributed computing. This paper is organized as follows. After introducing the TCM with a large number of artificial atoms in Sec. II, we introduce numerical method and its distributed computing transformation in Sec. III. We present the results of our numerical simulations in Sec. IV.1 and study the efficiency of distributed computing method in Sec. IV.2. Some brief comments on our results in Sec. V close out the paper.
@@
II Tavis–Cummings model with a large number of artificial atoms
@@ III.1 Numerical method
In order to apply distributed computing algorithm, the matrix exponential parts in the Eq. (6) needs to be converted into matrix multiplication-addition operations. Here we use the Taylor series approximation
@@ IV.1 The interaction between photons and matter
Figure 5: (online color) Comparison of time cost.

### 06 Apr 2024 v2

@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology [1, 2, 3]. When simulating chemical and biological reactions, a large number of particles are often involved, and the dimension of the quantum system composed of these particles increase exponentially as the number of particles increases, thus causing a problem called the curse of dimensionality — raised by R.E. Bellman [4, 5] to describe a number of events that occur while organizing and evaluating data in high-dimensional areas. The curse of dimensionality has always been a major obstacle to the study of high-dimensional quantum systems. However, the study of the structure of biological macromolecules is one of the most important frontier studies of quantum computing, thus the curse of dimensionality is an urgent issue to be solved. With the development of supercomputers in recent decades, the use of distributed computing can solve a series of memory and efficiency problems caused by the curse of dimensionality to a certain extent. Moreover, some distributed computing algorithms can be applied to quantum unitary evolution. A key contribution of this paper is the cavity quantum electrodynamics (QED) model, which is easy to implement in the laboratory and offers a unique scientific paradigm for studying light-matter interaction. According to this paradigm, impurity two-level systems, also known as atoms, are connected to fields of cavities. The cavity QED model includes the Jaynes–Cummings model (JCM) [6] and the Tavis-Cummings model (TCM) [7] as well as their generalizations. Many studies have been conducted recently in the field of these models [8, 9, 10, 11, 12]. In our previous studies, some methods have been used to solve some of the computing problems caused by the curse of dimensionality [13, 14]. In this paper, we try to simulate the cavity QED model with a large number of two-level artificial atoms on a supercomputing platform to reveal the rules of unitary evolution of a complex quantum system, and study the time cost and speedup under different distributed computing strategies comparing with that under the situation without distributed computing. This paper is organized as follows. After introducing the TCM with a large number of artificial atoms in Sec. II, we introduce numerical method and its distributed computing transformation in Sec. III. We present the results of our numerical simulations in Sec. IV.1 and study the efficiency of distributed computing method in Sec. IV.2. Some brief comments on our results in Sec. V close out the paper.
II Tavis–Cummings model with a large number of artificial atoms
@@ III.1 Numerical method
In order to apply distributed computing algorithm, the matrix exponential parts in the Eq. (6) needs to be converted into matrix multiplication-addition operations. In the past few decades, many calculation methods for exponential matrices have been proposed [16, 17]. Here we use the Taylor series approximation
@@ IV.1 The interaction between photons and matter
Figure 5: (online color) Comparison of time cost. (a) shows the comparison of time cost of distributed computing Taylor series approximation, (b) shows the comparison of time cost of distributed computing unitary evolution.
[…]
@@ IV.2 Comparison of cost time and speedup of different distributed computing strategies
Figure 6: (online color) Comparison of speedup. (a) shows the comparison of speedup of distributed computing Taylor series approximation, (b) shows the comparison of speedup of distributed computing unitary evolution.